OpenStudy (anonymous):

CAN SOME 1 HELP ME WITH SLOPE

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what you wanna know?

OpenStudy (anonymous):

THE EQUATION AND BASIC NECESSITIES

OpenStudy (anonymous):

it is really easy when you know formula

OpenStudy (brrandyn):

Slope is defined as: \[\frac{ Y_{2} - Y_{1} }{ X_{2}-X_{1}}\]

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] You already asked about this and i answered.

OpenStudy (anonymous):

formula is y2-y1/x2-x1

OpenStudy (alexandervonhumboldt2):

use slope formuaa\[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (anonymous):

can you help me

OpenStudy (alexandervonhumboldt2):

sure

OpenStudy (alexandervonhumboldt2):

xD

OpenStudy (anonymous):

alexander von i always wodered ur a preofessor at a college or smthing like that

OpenStudy (anonymous):

you have a graph for example you know that x=0 another x (it will be x2)=4 y=6 y2=12 and you put numbers in formula 12-6/4-0=3